Week of 150309

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Monday

Attendance:

  • local: Stefan (SCOD), Andrew (LHCb), Maarten (ALICE), Alessandro (ATLAS), Lorena (Databases), Ben (Grid Services), Herve (Storage),
  • remote: Hung-Te (ASGC), John (RAL), Michael (BNL), Sang-Un (KISTI), Rolf (IN2P3), Jose (CMS), Dmitri (KIT), Pepe (PIC), Di Qing (TRIUMF), Onno (NL-T1), Kyle (OSG),

Experiments round table:

  • ATLAS
    • Central Services
      • VOMS Admin GGUS:112177 . solution found (upgrade to java 8) need to be tested.
      • Request FTS3 update of package fts-rest-3.2.32.1-1.el6.noarch. This gives a performance increase for Rucio polling. Non-urgent and no downtime required.

  • CMS
    • Nothing to report. Very quiet weekend.

  • ALICE -
    • high activity

  • LHCb
    • Distributed computing dominated by Monte Carlo and user activities.
    • T0: SRM downtime completed ok for us.
    • T1:
      • CNAF and RRCKI ARCHIVE (tape) currently banned until nearline SRM endpoints working.

Sites / Services round table:

  • ASGC: Downtime on Friday 2am-4am UTC for upgrade of DPM servers
  • BNL: NTR
  • CNAF: NR
  • FNAL: NR
  • GridPP: NR
  • IN2P3: Last Thu LHCb mentioned problem with SE, didnít find any problem. Ask LHCb to file ticket.
  • JINR: NR
  • KISTI: Since last Friday failure of ALICE/VoBox due to connection failures to CERN/VOMS servers, caused by the primary link down, secondary routing was automatically configured but not working correctly. Checked this morning, only solution for secondary path is to connect to OPN needs to be discussed with local network experts.
  • KIT: NTR
  • NDGF: NR
  • NL-T1: dCache upgrade to 2.10 in progress, proceeding as expected.
  • NRC-KI: NR
  • OSG: NTR
  • PIC: Tomorrow DT for network upgrade & patch to dCache instance between 10.00 to 16.00 local time.
  • RAL: NTR
  • RRC-KI: NR
  • TRIUMF: NTR

  • CERN batch and grid services: Interruption for 1 hour on Friday in FTS due to increase in logging, issue is resolved and cause is being investigated
  • CERN storage services: NTR
  • Databases: CMS DB patches this week: CMSADG (tomorrow) and CMSARC (Wed)
  • GGUS: NR
  • Grid Monitoring: NR
  • MW Officer: NR

AOB:

Thursday

Attendance:

  • local: Stefan (SCOD), Andrew (LHCb), Lorena (Databases), Pablo (Monitoring&GGUS), Andrew (MW), Maarten (ALICE),
  • remote: Jens (NDGF), Lisa (FNAL), Denis (NL-T1), Sang-Un (KISTI), Thomas (KIT), David (ATLAS), Kyle (OSG), Di Qing (TRIUMF), Hung-Te (ASGC), Rolf (IN2P3), Pepe (PIC), Gareth (RAL)

Experiments round table:

  • ATLAS
    • Nothing special to report.

  • CMS
    • Nothing to report. No significant problems since Monday.

  • ALICE -
    • NTR

  • LHCb
    • Distributed computing dominated by Monte Carlo and user activities.
    • T0: VOMS tickets (GGUS:112281, GGUS:112279)
    • T1:
      • Tier1s back from storage downtimes, thank you.

Sites / Services round table:

  • ASGC: NTR
  • BNL: NR
  • CNAF: NR
  • FNAL: NTR
  • GridPP: NR
  • IN2P3: Urgent outage (announced) on Tuesday for correction of dCache
  • JINR: NR
  • KISTI: NTR
  • KIT: NTR
  • NDGF: NTR
  • NL-T1: NTR
  • OSG: NTR
  • PIC: NTR
  • RAL: Scheduled warning in GOCDB this morning, couple of 10 minute breaks. OUTAGE declared for Storage/CASTOR on April 8th
  • TRIUMF: NTR

AOB:

Edit | Attach | Watch | Print version | History: r13 < r12 < r11 < r10 < r9 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r13 - 2015-03-12 - PabloSaiz
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback